OpenStudy (anonymous):

Solve. –35 = 7t -1/5 –5 28 42

OpenStudy (anonymous):

i figured B but then i seen the negative so i guessed A

OpenStudy (texaschic101):

-35 = 7t (divide both sides by 7) -35/7 = 7t/7 -5 = t

OpenStudy (anonymous):

so it is B

OpenStudy (texaschic101):

yes..it is b -35 = 7 * -5 -35 = -35 (correct)